OpenStudy (anonymous):

Write an expression in simplest form for the area of a triangle with a base of 9y and a height of 4xy.

OpenStudy (vishweshshrimali5):

area = 1/2 * 9y * 4xy = 18xy^2

OpenStudy (anonymous):

area of a triangle = 1/2 * base * height base = 9y height = 4xy \[A= \frac{1}2*9y*4xy\] \[A=9y*2xy\] \[A=18xy^2\]